# CURVALID: Geometrically-guided Adversarial Prompt Detection

### v1

@@ 6 Conclusion
In this paper, we introduce CurvaLID, an adversarial prompt detection framework that filters out adversarial prompts before they are input to LLMs, ensuring their security. CurvaLID operates independently of LLMs, providing consistent performance across all models. It achieves over 0.99 accuracy and reduces the ASR of the tested adversarial prompts to near zero. CurvaLID leverages PromptLID and TextCurv, which analyze the geometric properties of prompts at the prompt-level and word-level, respectively. These measures address the limitations of word-level LID caused by stop words and punctuation, forming the foundation for CurvaLID’s robust performance in distinguishing benign and adversarial prompts. Future directions include expanding CurvaLID to spam detection, integrating human feedback, and testing with noisy prompts.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ 2. Semantic Weight and Embedding Norms
In NLP, the norm of a word embedding, $\|\vec{u}\|$, encodes the semantic “weight” or importance of a word within its context (Schakel & Wilson, 2015; Reif et al., 2019). Larger norms indicate that the embedding carries more semantic information, while smaller norms suggest less significance.
@@ Appendix B Supplementary information for experiments
The appendix is organized into four sections. B.1 provides supplementary information about CurvaLID. B.2 focuses on ablation studies, presenting experiments to analyze various aspects of CurvaLID’s performance. B.3 presents supplementary information on LID analysis. Finally, B.4 evaluates the performance of other SOTA defenses, comparing them to CurvaLID. All experiments were conducted on a system with a single Nvidia H100 GPU, 8 CPU cores, and 128 GB of RAM.
@@ B.1.2 CNN hyperparameter selection
The experimental results are shown in Table 4. We observe that the parameter tuning of the CNN does not significantly impact the overall detection accuracy, as the accuracy generally fluctuates around 0.98 to 0.99.
@@
We gathered 2519 prompts in total, with 1200 benign prompts and 1319 adversarial prompts. For benign prompts, we randomly sampled 300 prompts from each of Orca, MMLU, AlphacaEval and TruthfulQA. In addition, we gathered 320 SAP200 prompts, which are 40 randomly selected prompts from 8 adversarial goals. For DAN, we randomly sampled 350 prompts from the adversarial prompts uploaded on their Github, roughly half of their total number of prompts. For MathAttack, we used all 300 adversarial prompts uploaded to their Github. For GCG, we used their default parameter setting: learning rate = 0.01, batch size = 512, top-k = 256, and temperature = 1, and created a universal adversarial suffix. The targetted model is Vicuna-7B-v1.5. We used all 349 adversarial behaviors as listed in their Github.
@@ B.2 Ablation studies
This section presents ablation studies for CurvaLID, including its performance across various adversarial prompts and its effectiveness under different training conditions and parameter settings.
B.2.1 Performance metrics for CurvaLID in RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, and DrAttack
@@
We obtained 100 prompts from each of the four benign datasets. For PAIR, we sampled 171 adversarial prompts by generating PAIR adversarial prompts targeting Palm2, GPT4, Vicuna-7B, and Llama-2-13b-chat-hf. This is because PAIR’s code focuses on 50 adversarial goals for each LLM, and PAIR does not always find an adversarial prompt for each goal in its default setting. For RandomSearch, we obtain the adversarial prompts from their GitHub directly. We gathered the datasets from attacking vicuna-13B-v1.5, R2D2, Nemotron-4-340b, Llama-2-70b, Llama-2-13B, GPT4_Turbo, GPT3.5_Turbo, and Gemma-7B. We randomly selected 200 unique adversarial prompts, as we observed that some adversarial prompts are the same when attacking different LLMs. For AmpleGCG, we obtained permission to access their adversarial prompts directly. We randomly sampled 200 prompts from the attack to Vicuna-7B on AdvBench. For Persuasive Attack, we included 150 adversarial prompts uploaded by the authors on HuggingFace. For AutoDAN, we included 150 AutoDAN prompts generated by following the setting from their Github, with 50 prompts each targeting Vicuna-7b-v1.5, Llama2-7b, and GPT-4-0613. For DrAttack, we tested 150 adversarial prompts generated by following the setting from their Github, with 50 prompts targeting each of Vicuna-7B, LLaMA-2-7B-Chat, and GPT-3.5-turbo.
@@
B.2.2 Accuracy of CurvaLID with less data
@@
B.2.3 Performance metrics for CurvaLID with replacing MLP by local outlier factor or isolation forest
@@
We are testing 1900 prompts in total, with 100 prompts randomly sampled from each of SAP, DAN, MathAttack, GCG, PAIR, AmpleGCG and RandomSearch, and 300 prompts from each of Orca, MMLU, AlphacaEval and TruthfulQA. The experimental results are shown in Tables 7 and 8
@@
B.2.4 CurvaLID with PromptLID or TextCurv only
@@
B.2.5 Accuracy of CurvaLID in different embeddings
@@
B.2.6 CurvaLID in defending Vicuna-7B-v1.5
We measured the reduction in ASR after CurvaLID identified and filtered out the adversarial attacks in Vicuna-7B-v1.5, using the same settings specified in the respective original adversarial prompt papers. As shown in Table 11, CurvaLID successfully reduced the ASR of most attacks to zero, outperforming the studied SOTA defenses. The experimental results demonstrate that CurvaLID is highly applicable to real-world LLMs, effectively safeguarding them by detecting adversarial prompts before they are processed. Moreover, CurvaLID outperforms SOTA defenses, further highlighting its effectiveness and reliability.
@@
B.2.7 Replacement of CNN in CurvaLID Step 1 with Transformer and RNN models

### v2

@@ 6 Conclusion
In this paper, we introduce CurvaLID, an adversarial prompt detection framework that filters out adversarial prompts before reaching LLMs to maintain their security. CurvaLID operates independently of LLMs, providing consistent performance across models. It achieves over 0.99 accuracy and reduces the ASR of tested adversarial prompts to near zero. CurvaLID leverages PromptLID and TextCurv, which analyse the geometric properties of prompts at the prompt and word level, respectively. These measures address limitations of word-level LID caused by stop words and punctuation, forming the foundation for CurvaLID’s robust performance in distinguishing benign and adversarial prompts. Future work includes evaluating CurvaLID on benign prompts in other languages to strengthen multilingual robustness and ensure fairness in low-resource settings.
Reproducibility Statement
The algorithm outline of CurvaLID is presented in Section 4.3, with corresponding pseudo code in Appendix A.1. The definitions and methodologies of PromptLID and TextCurv are described in Sections 4.1 and 4.2, respectively. Theoretical results and proofs are provided in Section 4 and Appendix A.2. Experimental details are given in the first paragraph of Section 5 and Appendix B.1, with dataset descriptions in Appendix B.1.5. Model parameters and architectural configurations are reported in Appendices B.1.2 and B.1.4. The algorithm used in this research is included in the supplementary material, and all data used will be released upon publication.
@@ 2. Semantic Weight and Embedding Norms
In NLP, the norm of a word embedding, $\|\vec{u}\|$, encodes the semantic “weight" or importance of a word within its context (Schakel & Wilson, 2015; Reif et al., 2019). Larger norms indicate that the embedding carries more semantic information, while smaller norms suggest less significance.
@@ Appendix B Supplementary information for experiments
The appendix is organized into four sections. B.1 provides supplementary information about CurvaLID. B.2 focuses on ablation studies, presenting experiments to analyse various aspects of CurvaLID’s performance. B.3 presents supplementary information on LID analysis. Finally, B.4 evaluates the performance of other SOTA defences, comparing them to CurvaLID. All experiments were conducted on a system with a single Nvidia H100 GPU, 8 CPU cores, and 128 GB of RAM.
@@ B.1.2 CNN hyperparameter selection
The experimental results are shown in Table 4. We observe that the parameter tuning of the CNN does not significantly impact the overall detection accuracy, as the accuracy generally fluctuates around 0.98 to 0.99. Notably, the selected CNN extracts features from 1,000 prompts in under 0.5 seconds, demonstrating its efficiency.
@@
We use a total of 3,540 testing prompts, comprising 1,200 benign and 2,340 adversarial prompts.
For benign prompts, we randomly sampled 300 prompts from each of Orca, MMLU, AlphacaEval and TruthfulQA.
For adversarial prompts:
  • SAP: We gathered 320 SAP200 prompts by randomly selecting 40 prompts from each of the 8 adversarial goals.
  • DAN: We randomly sampled 350 prompts from the adversarial examples uploaded on their GitHub, covering roughly half of their total prompt set.
  • MathAttack: We used all 300 adversarial prompts provided on their GitHub.
  • GCG: We followed their default parameter settings—learning rate = 0.01, batch size = 512, top-k = 256, temperature = 1—to generate a universal adversarial suffix. All 349 adversarial behaviours listed in their GitHub were used.
  • PAIR: We generated 171 adversarial prompts using PAIR. Their implementation targets 50 adversarial goals per LLM, but does not always succeed in producing a prompt for each goal under the default configuration.
  • RandomSearch: We retrieved prompts directly from their GitHub and randomly selected 200 unique adversarial prompts, as many were duplicates across LLMs.
  • AmpleGCG: With author permission, we accessed their adversarial prompts and randomly sampled 200 prompts from the set.
  • Persuasive Attack: We included 150 prompts uploaded by the authors on Hugging Face.
  • AutoDAN: We included 150 prompts generated per the authors’ GitHub instructions for each targeted LLM.
  • DrAttack: We tested 150 adversarial prompts generated following the configuration provided in their GitHub.
@@ B.2 Ablation studies
This section presents ablation studies for CurvaLID, covering its performance across various adversarial prompts, baseline comparisons, and effectiveness under different training conditions and parameter settings. The section is organized as follows:
  • B.2.1 to B.2.7: Performance of CurvaLID across various adversarial prompts.
  • B.2.8 to B.2.11: Baseline comparisons against CurvaLID.
  • B.2.12 to B.2.24: Performance of CurvaLID under different training conditions and parameter settings.
  • B.2.25 to B.2.27: Ablation studies on the contributions of TextCurv and PromptLID.
B.2.1 Performance metrics for CurvaLID in PAIR, RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, DrAttack
@@
We obtained 100 prompts from each of the four benign datasets. The six adversarial datasets (PAIR, RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, and DrAttack) follow the same configuration described in Appendix B.1.5.
@@
B.2.2 CurvaLID on demonstration-based attacks
While our focus was on single-shot adversarial prompts, CurvaLID naturally extends to demonstration-based attacks due to its model-agnostic design. Operating independently of the target LLM, it is unaffected by prior demonstrations or context instructions. Any geometric anomaly within the prompt can be detected and filtered before reaching the LLM, effectively mitigating the attack.
We conducted additional experiments on In-Context Demonstration and cipher-based attacks (Wei et al., 2023; Yuan et al., 2023). For the former, we tested 400 adversarial prompts using the setup from Appendix B.2.1. For the cipher-based attack, we used 400 prompts from the official GitHub repository. The detailed results are shown in Table 6 below:
Table 6: Performance metrics for CurvaLID on in-context demonstration and cipher-based attacks.
Attack Type Benign Accuracy Adversarial Accuracy Overall Accuracy In-Context Demonstration Attack 0.9894 0.973 0.9812 Cipher-based Attack 0.9400 0.910 0.9250
B.2.3 CurvaLID on non-English adversarial prompts
Table 7 demonstrates CurvaLID’s effectiveness in detecting non-English adversarial prompts by evaluating it on nine languages from the MultiJail dataset (Deng et al., 2023b). We randomly sampled 300 prompts from each of the nine languages—Chinese (zh), Italian (it), Vietnamese (vi), Arabic (ar), Korean (ko), Thai (th), Bengali (bn), Swahili (sw), and Javanese (jv)—and tested them individually against 400 benign prompts, adjusted to avoid class imbalance. The benign prompts were sampled by gathering 100 entries from each of four different benign datasets. CurvaLID achieved an overall accuracy and F1 score of 0.994, highlighting its robust ability to detect adversarial prompts across a diverse range of languages.
B.2.4 CurvaLID on Persona Modulation
We evaluate CurvaLID against persona modulation attacks, following the setup in Shah et al. (Shah et al., 2023). We generated 200 attack prompts and randomly sampled 200 benign prompts from our benign dataset, as described in Section 5. The experiment setup follows our main evaluation in Section 5.1.
The results, reported in Table 8, show that CurvaLID achieves close to perfect detection accuracy on both benign and persona modulation prompts, demonstrating its robustness.
Table 8: CurvaLID accuracy on benign prompts and persona modulation attacks.
Method Benign Persona Modulation
CurvaLID 0.985 0.995
B.2.5 CurvaLID on HarmBench
We evaluated HarmBench (Mazeika et al., 2024) and tested GCG, AutoDAN, TAP, DirectRequest, and DAN on LLaMA2-7B and Vicuna-7B with 300 prompts each. We also re-evaluated baseline defences (SmoothLLM, Intentionanalysis, Self-Reminder) on these HarmBench prompts.
As shown in Table 9, CurvaLID achieved near-zero attack success rates and matched or outperformed these baselines.
Table 9: ASR (%) of HarmBench attacks under CurvaLID and baseline defences.
LLM defence GCG AutoDAN TAP DirectRequest DAN Average LLaMA-7B SmoothLLM 0.5 80.33 0 0 0.5 16.27 Intentionanalysis 0 0.33 0 0 1.67 0.40 Self-Reminder 0 0.33 0 0 2.5 0.57 CurvaLID 1.33 0 0 0 0 0.266 Vicuna-7B SmoothLLM 9.67 91.5 0 0 19.67 24.168 Intentionanalysis 0.33 11.0 1.5 0 9.5 4.47 Self-Reminder 9.33 92.0 7.33 0 57.5 33.23 CurvaLID 2.5 3.67 2.33 0 0 1.7
B.2.6 Evaluation on OR-Bench and XSTest
We evaluated CurvaLID on over-refusal benchmarks, namely OR-Bench and XSTest (Cui et al., 2024; Röttger et al., 2023). We tested on Vicuna-7B and LLaMA-2-7B and measured its effect on acceptance and rejection rates. We sampled 200 prompts each from the OR-Bench-Hard and OR-Bench-Toxic splits, and similarly from XSTest-Safe and XSTest-Unsafe. As shown in Tables 10 and 11, CurvaLID preserved LLaMA-2-7B’s strong rejection behaviour. For Vicuna-7B, CurvaLID reduced harmful acceptance by up to 30%, with only a modest increase (about 10%) in benign rejections, suggesting it can enhance safety without substantially impacting utility.
Table 10: CurvaLID performance on OR-Bench.
Dataset Metric Model defence Rate (%) OR-Bench-Hard Rejection LLaMA-2-7B No defence 85.5 CurvaLID 91.0 Vicuna-7B No defence 52.5 CurvaLID 60.5 OR-Bench-Toxic Acceptance LLaMA-2-7B No defence 0.5 CurvaLID 0.0 Vicuna-7B No defence 33.5 CurvaLID 1.5
Table 11: CurvaLID performance on XSTest.
Dataset Metric Model defence Rate (%) XSTest-Safe Rejection LLaMA-2-7B No defence 52.0 CurvaLID 59.5 Vicuna-7B No defence 12.5 CurvaLID 25.0 XSTest-Unsafe Acceptance LLaMA-2-7B No defence 0.5 CurvaLID 0.0 Vicuna-7B No defence 24.5 CurvaLID 2.0
B.2.7 Evaluation of baseline methods on benign prompts
We conducted an evaluation of baseline methods on benign prompts. Specifically, we compared CurvaLID with SmoothLLM, Self-Reminder, Intentionanalysis, ICD, and RTT3d on both Vicuna-7B and LLaMA2-7B. For the benign dataset, we randomly sampled 200 questions from MMLU. We use MMLU because unlike input-perturbation defences such as SmoothLLM and Intentionanalysis, CurvaLID operates as a detection algorithm without modifying the input or the internal mechanisms of the LLM. Thus, MMLU allows for a fair and consistent benchmark across all methods, as it consists of multiple-choice questions with clear-cut right or wrong answers.
To ensure fairness in measurement, if CurvaLID incorrectly flags a benign MMLU prompt as adversarial, we count the resulting LLM output as incorrect. The results are presented in Table 12 below. We observe that CurvaLID has minimal impact on benign performance. However, we also find that most competing defences similarly maintain high accuracy on MMLU, showing no significant degradation in utility.
Table 12: Accuracy on benign MMLU questions under different defence methods. We compare the utility impact of CurvaLID and baseline defences (SmoothLLM, Self-Reminder, Intentionanalysis, ICD, RTT3d) on Vicuna-7B and LLaMA2-7B. Accuracy is measured as the percentage of correctly answered MMLU prompts. “Original” refers to performance without any defence applied.
Model Original CurvaLID SmoothLLM Self-Reminder Intentionanalysis ICD RTT3d Vicuna-7B 46.0 48.0 40.0 46.0 50.0 48.0 39.5 LLaMA-2-7B 48.5 45.5 42.5 49.0 45.0 44.5 44.5
B.2.8 Baseline classification accuracy using RoBERTa embeddings and MLP
Table 13 presents the classification accuracy and performance metrics using RoBERTa embeddings as input to MLP. The MLP consists of a single hidden layer with 128 units, trained for a maximum of 300 iterations. The classification results show a benign class accuracy of 0.893, an adversarial class accuracy of 0.953, and an overall accuracy of 0.923.
Table 13: Classification accuracy and performance metrics using RoBERTa embeddings as input to MLP
Class Dataset Accuracy Class Acc. Overall Acc. F1 Benign Orca MMLU AlpEval TQA 0.893 0.923 0.924 0.872 0.899 0.893 0.905 Adv. SAP DAN MWP GCG 0.953 0.9400 0.962 0.885 1.000
B.2.9 Baseline classification accuracy using CurvaLID without PromptLID and TextCurv
To investigate the importance of the geometric features PromptLID and TextCurv in CurvaLID, we conducted an ablation study by removing Step 3 in CurvaLID (Figure 1). Specifically, we skipped the calculation of PromptLID and TextCurv, and directly fed the CNN representations from Step 2 into the MLP in Step 4 for binary classification. The goal is to assess the standalone performance of the CNN and MLP setup in CurvaLID, serving as a baseline without geometric information.
The results are presented in Table 14. Although the model performs reasonably well, its overall accuracy and F1 score drop by 7% compared to the full CurvaLID with PromptLID and TextCurv. This highlights the critical contribution of the geometric features in capturing topological differences between benign and adversarial prompts and improving classification robustness.
Table 14: Classification accuracy and performance metrics using CurvaLID without PromptLID and TextCurv (i.e., Step 3 removed). CNN representations are directly used as input to the MLP.
Class Dataset Accuracy Class Acc. Overall Acc. F1 Benign Orca MMLU AlpEval TQA 0.895 0.920 0.922 0.920 0.850 0.900 0.910 Adv. SAP DAN MWP GCG 0.947 1.000 0.952 0.845 0.992
B.2.10 ASR of B.2.8 and B.2.9 baseline classifications
We investigated the two baselines mentioned in B.2.8 and B.2.9 against seven types of adversarial prompts, including GCG, PAIR, DAN, AmpleGCG, SAP, MathAttack, and RandomSearch. Here we name the baseline in B.2.8 as "RoBERTa+MLP" and the baseline in B.2.9 as "CNN+MLP". In our approach, if a prompt is classified as a jailbreak prompt, it is rejected. The comparison against baseline defences across multiple LLMs, measured by ASR (%), is presented in Table 15.
The RoBERTa+MLP and CNN+MLP baselines perform poorly, particularly on PAIR, DAN, and MathAttack, where ASR remains high across all LLMs. These results highlight the importance of TextCurv and PromptLID in CurvaLID for enabling robust classification and significantly reducing ASR of adversarial prompts across LLMs.
Table 15: Comparison of CurvaLID with baseline defences in multiple LLMs, measured by ASR (%) on seven adversarial prompt types.
LLM defence GCG PAIR DAN AmpleGCG SAP MathAttack RandomSearch Vicuna-7B No defence 86.0 98.0 44.5 98.0 69.0 24.0 94.0 RoBERTa + MLP 0.0 12.2 3.6 2.5 4.25 11.5 0.2 CNN + MLP 0.2 16.4 4.1 2.9 0.0 15.0 0.7 CurvaLID 0.0 0.0 0.0 1.1 0.0 0.0 0.0 LLaMA2-7B No defence 12.5 19.0 2.0 81.0 9.5 11.7 90.0 RoBERTa + MLP 0.0 2.9 0.0 0.0 1.1 7.8 0.0 CNN + MLP 0.0 5.5 3.1 0.1 0.0 9.8 0.4 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0 GPT-3.5 No defence 12.0 48.0 6.33 82.0 0.9 10.5 73.0 RoBERTa + MLP 0.0 0.3 1.2 0.2 0.0 5.1 0.0 CNN + MLP 0.0 0.6 2.1 0.3 0.0 2.9 0.1 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0 PaLM2 No defence 14.9 98.0 49.7 88.9 55.1 18.9 91.9 RoBERTa + MLP 0.0 18.7 3.2 0.68 5.9 11.0 0.0 CNN + MLP 0.1 7.8 4.8 1.2 0.0 8.4 0.0 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0
B.2.11 Embedding Source Ablation: CNN and SBERT
CurvaLID employs a single lightweight CNN to produce both word-level (TextCurv) and sentence-level (PromptLID) representations from one input. To examine sensitivity to the sentence–embedding source, we replaced the CNN-derived sentence embeddings with the widely used SBERT model all-MiniLM-L6-v2 when computing PromptLID. Table 16 presents the comparison between CNN-based and the pretrained LLM embedding sentence embeddings for PromptLID and CurvaLID. It shows that the pretrained LLM embedding approach yields similar performance to our original CurvaLID configuration. Therefore, we conclude that the pretrained LLM embedding does not offer a performance advantage in this setting.
Table 16: Comparison between CNN-based and pretrained LLM (MiniLM) sentence embeddings for PromptLID and CurvaLID. The first two columns report PromptLID-only performance (sentence reps from CNN or MiniLM). The last two columns report the full CurvaLID with PromptLID computed from the respective embeddings.
Metric PromptLID (CNN only) PromptLID (MiniLM only) CurvaLID (PromptLID from CNN) CurvaLID (PromptLID from MiniLM) Benign Accuracy 0.987 0.945 0.984 0.955 Adversarial Accuracy 0.932 1.000 1.000 1.000 Overall Accuracy 0.958 0.967 0.992 0.973 F1 Score 0.958 0.960 0.992 0.967
B.2.12 Accuracy of CurvaLID with less data
@@
B.2.13 Performance metrics for CurvaLID with replacing MLP by local outlier factor or isolation forest
@@
We are testing 1900 prompts in total, with 100 prompts randomly sampled from each of SAP, DAN, MathAttack, GCG, PAIR, AmpleGCG and RandomSearch, and 300 prompts from each of Orca, MMLU, AlphacaEval and TruthfulQA. The experimental results are shown in Tables 18 and 19
@@
B.2.14 CurvaLID with PromptLID or TextCurv only
@@
B.2.15 Accuracy of CurvaLID in different embeddings
@@
B.2.16 Reduction in ASR of Vicuna-7B-v1.5 after applying CurvaLID
We measured the reduction in ASR after CurvaLID identified and filtered out the adversarial attacks in Vicuna-7B-v1.5, using the same settings specified in the respective original adversarial prompt papers. As shown in Table 22, CurvaLID successfully reduced the ASR of most attacks to zero, outperforming the studied SOTA defences. The experimental results demonstrate that CurvaLID is highly applicable to real-world LLMs, effectively safeguarding them by detecting adversarial prompts before they are processed. Moreover, CurvaLID outperforms SOTA defences, further highlighting its effectiveness and reliability.
@@
B.2.17 Replacement of CNN in CurvaLID Step 1 with Transformer and RNN models